Solution of differential equation p^ 2 +5p+6=0 is...

1
Expert's answer
2022-02-24T06:39:53-0500

Solution

If p = du/dx solving given equation we will get "p_{1,2}=-2.5\pm\sqrt{{2.5}^2-6}=-2.5\pm0.5"

So given differential equation  may be rewritten in the form

(du/dx + 2)( du/dx + 3)=0

du1/dx=-2 and du2/dx = -3  => u1 = -2x + C, u2 = -3x + D

So given differential equation  has two solutions (C and D are arbitrary constants)

u1(x) = -2x + C

u2(x) = -3x + D